# Rule-based Data Selection for Large Language Models

### v1

@@ A.6.1 Model training
A.6.2 Evaluation Benchmarks
@@
A.6.3 Number of selected rules
@@
A.6.4 Size of Sampled Data
@@
A.6.5 Rule Correlation of selected rules
Here we provide in Table 9 the rule indices (in the range $\{0,1,\dots,49\}$) for the rules selected by DPP and random selection. The fullest of all generated $50$ rules for each domain is provided in A.6.7. For each set of selected rules, we also calculate their rule correlation value $\rho$ (defined in 1). We confirm that indeed DPP selects rules with lower rule correlation than random selected rules.
@@ A.6.5 Rule Correlation of selected rules
A.6.6 GPT 10 uncorrelated rules
Another straightforward rule generation method is to directly prompt GPT-4 to generate 10 uncorrelated rules and rely on its understanding of the correlation between the rules. We have explored this by using a similar rule generation prompt as in A.6.7, where we provide the task description and data description, but this time we request for 10 rules and added one sentence “make sure the rules are uncorrelated” to further require the independence of the rules. The 10 “uncorrelated” GPT rules are provided in Table LABEL:tab:Appendix-EvalB-GPT10UncorrelatedRules-Code and LABEL:tab:Appendix-EvalB-GPT10UncorrelatedRules-Math below. Following this, we rated the data according to the 10 GPT rules and calculated the rule correlation $\rho$ of the score vectors. We tested for Code and Math domain and got $\rho_{Code}=0.65$ and $\rho_{Math}=0.56$, both are significantly higher than DPP correlation values in Tabel 9. For Code, even random 10 rules selected from a pool of 50 rules provide lower correlation than the 10 rules directly generated by GPT that are claimed to be “uncorrelated”. This shows that our two-step approach—first generating enough rules to ensure diversity, followed by employing DPP on the rating vectors to select rules—is superior and also more task-specific.
@@ A.6.6 GPT 10 uncorrelated rules
A.6.7 Prompts and Generated Rules
For brevity, we provide the templates for both the rule generation and rating prompts for the Math domain. To adapt these templates for other domains, replace terms specific to Math (such as “mathematical tasks” and “mathematical reasoning and analysis”) with relevant terminology from the desired domain.
@@ A.6.7 Prompts and Generated Rules
Generated 50 rules for each of four domains: Note that the IMDB rules here are used to select data for LLM training, whereas the IMDB rules in A.5.5 are used for data comparison in order to eventually calculate quality scores for the $50$ IMDB reviews. Although similar, they are not the same set of rules.

### v2

@@ A.6.1 Model training
A.6.2 More Baseline Methods
Here we add two more baseline methods: LESS (Xia et al., 2024): selecting data based on the estimated data influences, and DiverseEvol (Wu et al., 2023): an iterative sampling algorithm to ensure data diversity. It is important to note that DiverseEvol focuses solely on a single quality aspect: the diversity of data, while our method ensures diversity across multiple rating aspects. Another remark is that in the original papers, these methods were specifically used for instruction tuning data. We copy the three rule-based methods in our framework from 1 for comparison purposes. From the results below, we see these two methods, while being computationally expensive, are not showing good performance under our experiment settings.
Method ARC-Easy ARC-Challenge Winogrande MMLU SST-2 Average
LESS 59.8 25.2 52.9 25.6 49.3 42.5
DiverseEvol 59.7 24.7 53.3 25.6 49.2 42.5
All 50 Rules 60 25.3 53 26.1 49.3 42.7
Random 10 Rules 60 25.1 53.7 25.8 49.5 42.8
DPP 10 Rules 60 25.7 54 26.2 50.1 43.2
Table 8: General continued pre-training of Pythia-1B using 20K selected data samples from SlimPajama.
Method IMDB Medical
LESS 46.6 23.6 40.4 24 29.3
DiverseEvol 51.1 23.6 42.5 23 29.7
All 50 Rules 51 23.1 42.6 23 29.6
Random 10 Rules 51.7 24 41.2 23.5 29.6
DPP 10 Rules 53.5 24.6 43.3 26.8 31.6
Table 9: IMDB & Medical fine-tuning on Pythia-1B, each using 20K selected data samples from SlimPajama.
Method Math Code
LESS 41.5 40.4 33 38.3 41.4 43.5 43.9 45.3 43.5
DiverseEvol 41.2 38.5 35 38.2 38.4 43.6 42.8 47.8 43.1
All 50 Rules 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.8 35.2 39.3 48.5 41 46.6 48.1 46
DPP 10 Rules 43.7 40.6 38 40.8 50.5 44.2 46.9 52.7 48.6
Table 10: Math & Code fine-tuning on Llama3-8B, each using 20K selected data samples from SlimPajama.
A.6.3 Variance of Trials
Due to computational resource constraints, we were unable to perform multiple repetitions of all experiments. However, as mentioned in Section 5, we conducted 3 independent trials in four domains for Uniform Sampling and methods involving randomness in rule selections, including GPT-Uncorrelated, Random 10 Rules, and DPP 10 Rules (note that DPP sampling is also non-deterministic) to mitigate the effects of randomness, and we report their standard deviations here.
Method IMDB Medical
Uniform Sampling $43.9_{1.1}$ $23.0_{0.42}$ $42.1_{0.62}$ $22.5_{0.71}$ $28.9_{0.77}$
GPT-Uncorrelated $50.9_{0.71}$ $23.7_{0.11}$ $42_{0.3}$ $22.7_{0.76}$ $29.4_{0.2}$
Random 10 Rules $51.7_{0.21}$ $24_{0.42}$ $41.2_{1.2}$ $23.5_{0.41}$ $29.6_{0.58}$
DPP 10 Rules $53.5_{0.58}$ $24.6_{0.37}$ $43.3_{0.43}$ $26.8_{0.76}$ $31.6_{0.41}$
Table 11: Mean and standard deviation over 3 independent trials for the IMDB & Medical fine-tuning setting.
Method Math Code
Uniform Sampling $40.5_{0.35}$ $39.2_{0.56}$ $35_{0.2}$ $38.2_{0.30}$ $38.7_{1.27}$ $38.2_{1.6}$ $38.2_{0.42}$ $39.7_{1.2}$ $38.7_{0.94}$
GPT-Uncorrelated $41.4_{0.17}$ $39_{0.21}$ $37.3_{0.57}$ $39.2_{0.28}$ $41.2_{0.15}$ $43.5_{0.11}$ $39.6_{1.44}$ $48.6_{0.15}$ $43.2_{0.32}$
Random 10 Rules $42.9_{0.1}$ $39.8_{1.3}$ $35.2_{0.28}$ $39.3_{0.48}$ $48.5_{0.6}$ $41_{0.94}$ $46.6_{0.85}$ $48.1_{1.2}$ $46_{0.78}$
DPP 10 Rules $43.7_{0.61}$ $40.6_{0.32}$ $38_{0}$ $40.8_{0.15}$ $50.5_{0.36}$ $44.2_{0.26}$ $46.9_{0.2}$ $52.7_{0.15}$ $48.6_{0.22}$
Table 12: Mean and standard deviation over 3 independent trials for the Math & Code fine-tuning setting.
Here we perform the $t$-test to demonstrate that the advantage of DPP 10 Rules is significant compared to other methods. We include the $t$-statistics and $p$-values in the table. If we choose the significance threshold $p=0.05$, then we see that all the comparisons are significant.
Comparison IMDB Medical average
DPP vs GPT-Uncorrelated $t=4.912$, $p=0.00881$ $t=8.353$, $p=0.00408$
DPP vs Uniform Sampling $t=13.371$, $p=0.00086$ $t=5.361$, $p=0.01218$
DPP vs Random 10 Rules $t=5.054$, $p=0.02255$ $t=4.877$, $p=0.01065$
Table 13: Comparison of DPP method with other methods in IMDB and Medical AVG domains using Welch’s t-test.
Comparison Math average Code average
DPP vs GPT-Uncorrelated $t=8.724$, $p=0.00293$ $t=24.085$, $p=0.00005$
DPP vs Uniform Sampling $t=13.426$, $p=0.00099$ $t=17.762$, $p=0.00195$
DPP vs Random 10 Rules $t=5.166$, $p=0.02415$ $t=5.557$, $p=0.02204$
Table 14: Comparison of DPP method with other methods in Math and Code domains using Welch’s t-test.
A.6.4 Evaluation Benchmarks
@@
A.6.5 Number of selected rules
@@
A.6.6 Size of Sampled Data
@@
A.6.7 Distribution of Selected Data
Evaluating and contrasting the quality of data subsets selected by different methods is challenging and often necessitates extensive human intervention. To address this, the authors examined the initial 100 examples selected by each method. This examination revealed notable distinctions in the relevance and domain specificity of the data selected. Specifically, our DPP rule-based approach demonstrated a marked ability to identify and select examples that were highly pertinent to specific domains. For instance, in experiments focused on the Code domain, this method favored the inclusion of data containing code. In contrast, other less targeted methods, such as QuRating and Uniform Sampling, often yield selections that lack domain-specific relevance. This insight underscores the efficacy of using tailored, rule-based methods over generic ones for tasks where domain alignment is critical.
Although it is hard to compare the distribution of the selected data, we provide a visual representation in Figure 10 below, showcasing the meta-data (categories of the data samples) distributions for the Code domain as a representative example. Notably, the DPP methods with 10 and 50 rules tend to select more data from GitHub and StackExchange for Code fine-tuning.
Moreover for IMDB domain, in Figure 11 we investigated the text length distribution. We see that the QuRating is very close to the original SlimiPajama distribution, where we conjecture that in this case the data distribution is very close to uniformly sampled data. The methods within our framework have a tendency toward longer texts. Additionally, in Figure 12 we use bigram entropy (the Shannon entropy of the distribution over the unique bigrams) as an indicator of the text diversity. We again see that the entropy distribution of QuRating is very close to the original SlimPajama, where our methods generally select data with higher entropy/diversity and the entropy distributions are more concentrated.
(a) DPP 10 rules
(b) Random 10 rules
(c) All 50 rules
(d) QuRating
(e) SlimPajama1M
Figure 10: Comparison of meta-data distribution across different methods. The last is the original distribution of our source data.
(a) DPP 10 rules
(b) Random 10 rules
(c) All 50 rules
(d) QuRating
(e) SlimPajama1M
Figure 11: Comparison of text length distribution across different methods. The last is the original distribution of our source data.
(a) DPP 10 rules
(b) Random 10 rules
(c) All 50 rules
(d) QuRating
(e) SlimPajama1M
Figure 12: Comparison of text diversity distribution across different methods. The last is the original distribution of our source data.
A.6.8 Rule Correlation of selected rules
Here we provide in Table 17 the rule indices (in the range $\{0,1,\dots,49\}$) for the rules selected by DPP and random selection (in one trial). The fullest of all generated $50$ rules for each domain is provided in A.6.11. For each set of selected rules, we also calculate their rule correlation value $\rho$ (defined in 1). We confirm that indeed DPP selects rules with lower rule correlation than random selected rules.
@@ A.6.8 Rule Correlation of selected rules
A.6.9 GPT 10 uncorrelated rules
Another straightforward rule generation method is to directly prompt GPT-4 to generate 10 uncorrelated rules and rely on its understanding of the correlation between the rules. We have explored this by using a similar rule generation prompt as in A.6.11, where we provide the task description and data description, but this time we request for 10 rules and added one sentence “make sure the rules are uncorrelated” to further require the independence of the rules. The 10 “uncorrelated” GPT rules are provided in Table LABEL:tab:Appendix-EvalB-GPT10UncorrelatedRules-Code and LABEL:tab:Appendix-EvalB-GPT10UncorrelatedRules-Math below. Following this, we rated the data according to the 10 GPT rules and calculated the rule correlation $\rho$ of the score vectors (in one GPT-Uncorrelated trial). We tested for Code and Math domain and got $\rho_{Code}=0.65$ and $\rho_{Math}=0.56$, both are significantly higher than DPP correlation values in Tabel 17. For Code, even random 10 rules selected from a pool of 50 rules provide lower correlation than the 10 rules directly generated by GPT that are claimed to be “uncorrelated”. This shows that our two-step approach—first generating enough rules to ensure diversity, followed by employing DPP on the rating vectors to select rules—is superior and also more task-specific.
@@ A.6.9 GPT 10 uncorrelated rules
A.6.10 Use GPT to select 10 uncorrelated rules
In this part, we discuss a very similar setting to the previous section. However, instead of directly prompting GPT to generate 10 rules, we let GPT to replace the role of DPP and select 10 “uncorrelated” rules out of the 50-rule pool. First, in Table 20 below, we calculate the rule correlation similarly as in Table 17. We see again although we prompt GPT-4 to select “uncorrelated” rules, the rule-correlation of the selected 10 rules are still higher than our DPP-selected rules in Table 17. Moreover, we fine-tuned with the selected data and benchmarked the LLM performance. From the results in Table 21 and Table 22, we again see that it underperforms compared to our method.
Domain Method Rule Indices
IMDB GPT selected 10 rules ($\rho=0.67$) [0, 1, 4, 10, 13, 17, 25, 31, 40, 49]
Medical GPT selected 10 rules ($\rho=0.40$) [0, 4, 7, 11, 15, 24, 29, 34, 42, 49]
Math GPT selected 10 rules ($\rho=0.56$) [0, 3, 7, 11, 17, 24, 28, 38, 44, 48]
Code GPT selected 10 rules ($\rho=0.65$) [0, 4, 9, 12, 16, 23, 29, 34, 43, 49]
Method IMDB Medical
GPT selected 10 rules 51.6 21.9 42.2 24 29.4
All 50 Rules 51 23.1 42.6 23 29.6
Random 10 Rules 51.7 24 41.2 23.5 29.6
DPP 10 Rules 53.3 24.3 43 26 31.1
Table 21: IMDB & Medical fine-tuning on Pythia-1B, each using 20K selected data samples from SlimPajama.
Method Math Code
GPT selected 10 rules 42.5 40.3 36 39.6 40.8 43.4 42.8 50.3 44.3
All 50 Rules 41.8 40.7 33 38.5 43.9 43.4 46.6 49.1 45.8
Random 10 Rules 42.9 39.6 35 39.2 48.5 40.8 46.6 48.1 46
DPP 10 Rules 43.6 40.4 38 40.7 50.6 44.1 46.9 52.8 48.6
Table 22: Math & Code fine-tuning on Llama3-8B, each using 20K selected data samples from SlimPajama.
A.6.11 Prompts and Generated Rules
For brevity, we provide the templates for both the rule generation and rating prompts for the Math domain. To adapt these templates for other domains, replace terms specific to Math (such as “mathematical tasks” and “mathematical reasoning and analysis”) with relevant terminology from the desired domain. We use GPT-4 to help us generate these task description and data descriptions.
@@ A.6.11 Prompts and Generated Rules
Generated 50 rules for each of four domains: Note that the IMDB rules here are used to select data for LLM training, whereas the IMDB rules in A.5.6 are used for data comparison in order to eventually calculate quality scores for the $50$ IMDB reviews. Although similar, they are not the same set of rules.